OpenStudy (anonymous):

PLEASE, can anyone help me with what to do? Soybean meal is 18% protein, cornmeal is 9% protein. How many pounds of each should be mixed together in order to get 360-lb mixture that is 17% protein? This makes no sense to me.

OpenStudy (anonymous):

ok let me write it out

OpenStudy (anonymous):

Oh Thank yo so much!!!

OpenStudy (anonymous):

call the amount of soybean x so the amount of cornmeal is \[360-x\] because they have to total 360 when you add them. soybean is 18% protein so you will have is \[.18\times x\] protein from the soybean and also \[.09\times (360-x)\] protein from the cornmeal. you want this to be \[.17\times 360\] so set them equal and solve \[.18x+.09\times (360-x)=.17\times 360=61.2\]

OpenStudy (anonymous):

i will be happy to show you how to solve this if you like. if the decimals are annoying multiply both sides by 100 to get \[18x+9(360-x)=6120\] \[18x+3240-9x=6120\] \[9x+3240=6120\] \[9x=6120-3240=2880\] \[x=\frac{2880}{9}=320\]

OpenStudy (anonymous):

if not clear let me know

OpenStudy (anonymous):

Thank you very much, the 320 is that the soybean or cornmeal? I am math dumb I'm sorry